OpenStudy (anonymous):

what is the radius of a circle with an equation: x^2 + y^2 + 2x + 4y -9 = 0

OpenStudy (sidsiddhartha):

do u know what is general form of a circle?

OpenStudy (sidsiddhartha):

its \[x ^{2}+y ^{2}+2gx+2fy+c=0\]

OpenStudy (dumbcow):

use completing the square --> (x^2 +2x) + (y^2 +4y) = 9 --> (x^2 +2x + 1) + (y^2 + 4y + 4) = 9 + 1+4 --> (x+1)^2 + (y+2)^2 = 14 r^2 = 14

OpenStudy (wwe123):

(x^2 +2x) + (y^2 +4y) = 9 (x^2 +2x + 1) + (y^2 + 4y + 4) = 9 + 1+4 (x+1)^2 + (y+2)^2 = 14 r^2 = 14

OpenStudy (sidsiddhartha):

in your problem g=1 and f=2 and c=-9 now rad\[\sqrt{g ^{2}+f ^{2}-c}=\sqrt{1+4+9}=\sqrt{14}\]ius=